Solve the initial value problem by using La Place:

y 2 y-2u(t - 2n) y(0) 0, y(0)= el-2T
Answer #1

Solution: a) y +2y+2y= et-20 (4-20) y (0=0, y colso taking Laplace transform Lay+2yltzy} = Lof et-zury-ang} Y(8)82-466)8-y- 8- 3 22 2 expanding and rearranging CSf) Yola peste enter teilcom 2 Laplace a inverse taking transform [f4c8l=é C C4-25) -